# Learning to Control Unknown Strongly Monotone Games

### v1

@@ 4.2 Finite Time Convergence Guarantees
In this subsection we detail how to obtain our finite time convergence guarantees. This result also uses the non-expansiveness on $g$, but requires a more nuanced argument. To prove the finite time mean square error bound, we bound the error as follows
@@
We first bound the mean square distance between $\boldsymbol{x}_{t}$ and $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, i.e., the NE corresponding to $\boldsymbol{\alpha}_{t}$.
@@
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that
@@
for all $t>T_{0}$.
The key property used in the proof for this lemma is the strongly monotone property of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$. This recursive bound gives a rate of $O(\eta_{t})$ when solved.
The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$. Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) Krasnosel’skiĭ–Mann algorithm. If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [64, 65]. Since this assumption is not possible in our case, we use a different approach and employ the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
@@
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that
@@
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}},$
for all $t>T_{0}$.
Lemma 7 and 8 together complete the proof for Theorem 2.
@@ 5 Simulation Results
In this section, we provide numerical simulations that include the two applications discussed in Section 2.1. For all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$. We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\left\{\alpha_{0}^{k}\right\}$ and $\left\{x_{n,0}^{k}\right\}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
For the resource allocation game (Figure 2), we design our game taking inspiration from DSM in electricity grid [66, 67]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as
$\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$
Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. The parameter $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$. We use the reward function
$r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$
Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{N}^{i}$, i.e., the weighted load at resource $i$. The coefficients $a_{n,i}^{(i)}$ are sampled independently from a Gaussian distribution with mean $1$ and variance $0.2$. The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are sampled independently from a Gaussian distribution with mean $35$ and variance $5$.
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a Nash equilibrium where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
@@
We work with $N=100$ players and $d=5$, and action sets $\sum_{i=1}^{d}x_{n}^{i}\leq 10$ for each player $n$. To generate a random $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a Gaussian distribution with mean 0 and variance 1. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained in a similar method as $\boldsymbol{G}$ but by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.

### v2

@@ 4.2 Finite Time Convergence Guarantees
In this subsection we detail how to obtain our finite time convergence guarantees. This result also uses the non-expansiveness on $g$, but requires a more nuanced argument that bounds the error as follows
@@
We first bound the mean square distance between $\boldsymbol{x}_{t}$ and $\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$, i.e., the NE corresponding to $\boldsymbol{\alpha}_{t}$. The key to proving this lemma is the strong monotonicity of the game which gives a recursive bound on $\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$. This recursive bound gives a rate of $O(\eta_{t})$ when solved.
@@
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime}>0$ and $T_{0}$ such that for all $t>T_{0}$:
$\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]\leq C^{\prime}\eta_{t}.$
Proof.
Returning to the the recursive bound (4.1) from the proof for Lemma 6, and using Lemma 6 that shows that $\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]\leq C_{0}$, we now obtain the recursion
$\mathbb{E}[\|\boldsymbol{x}_{t+1}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t+1})\|^{2}]$ $\leq(1-1.5\mu\eta_{t})\mathbb{E}[\|\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})\|^{2}]$
$+\eta_{t}^{2}(c_{4}+2c_{2}^{2}C_{0}).$
For $\eta_{t}=(t+T_{1})^{-\eta}$ with $10<\eta<$ 1, this recursion yields [12]
@@
for $t>T_{0}$ for some constant $C^{\prime}$. ∎
The next lemma bounds the mean squared distance between $\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})$ and $\boldsymbol{\ell^{*}}$. Note that we need to bound $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ as $g(\boldsymbol{\alpha})-\boldsymbol{\alpha}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha})-\boldsymbol{\ell^{*}})$. We prove the bound in the following lemma using Lemma 7 and a technique similar to the bound for (inexact) KM algorithm. If we could assume that $\epsilon_{t}>\epsilon^{*}$ for some $\epsilon^{*}$, then $\mathbb{E}[\|g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}\|^{2}]$ is known to decay at the rate $o(1/t)$ [63]. Since this assumption is not true in our case, we use the assumptions on the stepsize sequences to get a convergence rate of $O(1/(t\epsilon_{t}))$.
@@
Let Assumptions 1-5 be satisfied. Then, there exist $C^{\prime\prime}>0$ and $T_{0}>0$ such that for all $t>T_{0}$.
$\mathbb{E}[\|\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}}\|^{2}]\leq\frac{C^{\prime\prime}}{t\epsilon_{t}}.$
Proof.
Define $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}$, $q_{t}=\gamma\boldsymbol{A}(\boldsymbol{x}_{t}-\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t}))$ and $\epsilon_{t}^{\prime}=\epsilon_{t}/\gamma$. Returning to the recursive bound (4.1) from the proof of Lemma 6, and now using Lemma 6 and Lemma 7,
$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}-\epsilon_{t}^{\prime}(1-\epsilon_{t}^{\prime})\mathbb{E}[\|p_{t}+q_{t}\|^{2}]$ (17)
for some constant $c_{5}$. For the third term above, we note that
$-\|p_{t}+q_{t}\|^{2}$ $=-\|p_{t}\|^{2}-\|q_{t}\|^{2}-2\langle p_{t},q_{t}\rangle$
$\leq-\|p_{t}\|^{2}-\|q_{t}\|^{2}+2\|p_{t}\|\|q_{t}\|$
$\leq-\|p_{t}\|^{2}+c_{6}\|q_{t}\|.$
Note that $p_{t}=g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t}=\gamma(\boldsymbol{A}\boldsymbol{x^{*}}(\boldsymbol{\alpha}_{t})-\boldsymbol{\ell^{*}})$. Hence the last inequality follows from the compactness of $\mathcal{X}$ for some constant $c_{6}>0$. This implies $\mathbb{E}\left[-\|p_{t}+q_{t}\|^{2}\right]\leq\mathbb{E}\left[-\|p_{t}\|^{2}\right]+2c_{7}\sqrt{\eta_{t}},$ for some constant $c_{7}>0$. This inequality follows from Lemma 7 and since $\mathbb{E}[\|q_{t}\|]^{2}\leq\mathbb{E}[\|q_{t}\|^{2}]$.
Using our assumption that $\epsilon_{t}$ goes to zero, we have that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$. For the rest of the proof, we pick $T_{0}$ to be large enough such that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$, $\epsilon_{t}<\gamma/2$ and $\eta_{t}<C_{2}$ for all $t>T_{0}$, for appropriate $C_{1}$ and $C_{2}$. Returning to (4.2), for all $t>T_{0}$, we have
$\mathbb{E}\left[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}\right]$
@@
$\leq\mathbb{E}\left[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}\right]+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}-(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}].$
This gives us
$(\epsilon_{t}^{\prime}/2)\mathbb{E}[\|p_{t}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{t}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;\;+\eta_{t}^{2}c_{5}^{2}+\epsilon_{t}^{\prime}\sqrt{\eta_{t}}c_{7}.$
Summing from $i=T_{0}$ to $t$ gives us
$\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\mathbb{E}[\|p_{i}\|^{2}]$ $\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]-\mathbb{E}[\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7}).$ (18)
The left hand term in the above inequality depends on $\|p_{i}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$. We next show a bound for $\|p_{i}\|^{2}$ in terms of $\|p_{t}\|^{2}$ for all $i\in\{T_{0},\ldots,t\}$. For this, we first recall that $\boldsymbol{\alpha}_{t+1}=\boldsymbol{\alpha}_{t}+\epsilon_{t}^{\prime}(p_{t}+q_{t}).$ Next,
$\|p_{t+1}\|^{2}=\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+2\langle p_{t+1}-p_{t},p_{t}\rangle$
$\leq\|p_{t}\|^{2}+\|p_{t+1}-p_{t}\|^{2}+\frac{2}{\epsilon_{t}^{\prime}}\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}-\epsilon_{t}^{\prime}q_{t}\rangle$
Now,
$2\langle p_{t+1}-p_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t},\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle$
$=2\langle g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t}),\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\rangle-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}+\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-2\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}$
$\;\;-\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})-\boldsymbol{\alpha}_{t+1}+\boldsymbol{\alpha}_{t}\|^{2}$
$=\|g(\boldsymbol{\alpha}_{t+1})-g(\boldsymbol{\alpha}_{t})\|^{2}-\|\boldsymbol{\alpha}_{t+1}-\boldsymbol{\alpha}_{t}\|^{2}-\|p_{t+1}-p_{t}\|^{2}$
$\leq-\|p_{t+1}-p_{t}\|^{2}.$
Here the final inequality follows from the non-expansive nature of $g(\cdot)$. This gives us
$\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\|p_{t+1}-p_{t}\|^{2}-2\langle p_{t+1}-p_{t},q_{t}\rangle$
$=\|p_{t}\|^{2}-\frac{1-\epsilon_{t}^{\prime}}{\epsilon_{t}^{\prime}}\left\|p_{t+1}-p_{t}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}q_{t}\right\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}$
$\leq\|p_{t}\|^{2}+\frac{\epsilon_{t}^{\prime}}{1-\epsilon_{t}^{\prime}}\|q_{t}\|^{2}.$
Using the assumption that $20<\epsilon_{t}^{\prime}<1/$ 2 for all $t>T_{0}$, this gives us $\|p_{t+1}\|^{2}\leq\|p_{t}\|^{2}+2\epsilon_{t}^{\prime}\|q_{t}\|^{2}.$ Taking expectation and using Lemma 7, this results in the bound
$\mathbb{E}[\|p_{i}\|^{2}]$ $\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2\epsilon_{j}^{\prime}\mathbb{E}\|q_{j}\|^{2}$
$\geq\mathbb{E}[\|p_{t}\|^{2}]-\sum_{j=i}^{t-1}2c_{8}\epsilon_{j}^{\prime}\eta_{j},$
for some constant $c_{8}>0$. Returning to (4.2), we now have
$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq\mathbb{E}[\|\boldsymbol{\alpha}_{T_{0}}-\boldsymbol{\alpha^{*}}\|^{2}]$
$\;\;+\sum_{i=T_{0}}^{t}(\eta_{i}^{2}c_{5}^{2}+\epsilon_{i}^{\prime}\sqrt{\eta_{i}}c_{7})+2c_{8}\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}.$ (19)
Now we use our assumption that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ for all $t>T_{0}$. Then $\epsilon_{i}^{\prime}\sqrt{\eta_{i}}=O(\eta_{i}^{2})$ and hence the sequence $\{\epsilon_{i}^{\prime}\sqrt{\eta_{i}}\}$ is summable. Finally for the double summation, we use the integral test and Assumption 5 to obtain
$\sum_{i=T_{0}}^{t}\epsilon_{i}^{\prime}\sum_{j=i}^{t-1}\epsilon_{j}^{\prime}\eta_{j}\leq\frac{1}{\gamma^{2}}\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\sum_{j=i}^{\infty}\frac{1}{(j+T_{3})^{\epsilon+\eta}}$
$\leq\sum_{i=T_{0}}^{\infty}\frac{1}{(i+T_{2})^{\epsilon}}\frac{c_{9}}{(i+T_{3})^{\epsilon+\eta-1}}\leq\sum_{i=T_{0}}^{\infty}\frac{c_{10}}{(i+T_{3})^{4\eta-1}}.$
Here $c_{9}$ and $c_{10}$ are constants and $T_{3}=\max\{T_{1},T_{2}\}$. The final inequality follows from the fact that $\epsilon_{t}^{2}\leq C_{1}\eta_{t}^{3}$ and hence $2\epsilon\geq 3\eta$. Since $\eta_{t}$ is square summable, $\eta>0.5$ and hence $14\eta-1>$ 1, so the summation is bounded by a constant. Hence all terms in the right-hand side of (4.2) are bounded by a constant. Then for some $c_{11}>0$, we have
$\mathbb{E}[\|p_{t}\|^{2}]\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\leq c_{11}.$
Now $\sum_{i=T_{0}}^{t}(\epsilon_{i}^{\prime}/2)\geq(t-T_{0}+1)\epsilon_{t}^{\prime}/2$. This finally gives us the bound $\mathbb{E}[\|p_{t}\|^{2}]\leq C^{\prime\prime}/(t\epsilon_{t}),$ for all $t>T_{0}$. ∎
@@ 5 Simulation Results
In this section, we simulate the two applications discussed in Section 2.1. In all experiments, we use the step-size sequence $\eta_{t}=\frac{1}{\left(t+1\right)^{0.5}}$ and the control step-size sequence $\varepsilon_{t}=\frac{1}{\left(t+1\right)^{0.75}}$. We add Gaussian noise with mean 00 and variance $\sigma^{2}=\frac{1}{4}$ to the gradients of each player. We ran 100 realizations for each experiment and plotted the average result along with the standard deviation region, which was always small. In each realization, $\alpha_{0}^{k}$ and $x_{n,0}^{k}$ were chosen uniformly and independently at random on $\left[0,2\right]$ and $\left[0,0.1\right]$, respectively.
We pick a resource allocation game (Figure 2) inspired by DSM in electricity grid [64]. We keep $N=1000$ and $d=24$. We choose the action set for each player $n$ as $\mathcal{X}_{n}=\{\boldsymbol{x}_{n}|0\leq x_{n}^{i}\leq\mu_{n}^{i},\sum_{i=1}^{d}x_{n}^{i}\leq\mu_{n}\}.$ Here $\mu_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$, and $\mu_{n}$ was sampled independently for each $n$ from a Gaussian distribution with mean $5$ and variance $1$. We use the reward function $r_{n}(\boldsymbol{x})=\sum_{i=1}^{d}\omega_{n}^{i}x_{n}^{i}-0.3(x_{n}^{i})^{2}-0.01x_{n}^{i}s_{i}^{2}.$ Here $\omega_{n}^{i}$ was sampled independently for each $n,i$ from a Gaussian distribution with mean $1.5$ and variance $0.5$. We also defined $s_{i}=\sum_{n=1}^{N}a_{n,i}^{(i)}x_{n}^{i}$, i.e., the weighted load at resource $i$. The coefficients $a_{n,i}^{(i)}$ are i.i.d. samples from a Gaussian distribution with mean $1$ and variance $0.2$. The objective is $s_{i}=\ell_{i}^{*}$ where $\ell_{i}^{*}$ are i.i.d. samples from a Gaussian distribution with mean $35$ and variance $5$.
In Fig. 2(a), we plot the weighted load for each resource along with their target loads. We can observe that the system converges to a NE where the loads are balanced. In Fig. 2(b), we plot the weighted load for the first 3 resources. To study the effectiveness of our algorithm, we also plot the loads in the case of a uncontrolled system., i.e., where $\boldsymbol{\alpha}=\boldsymbol{0}$. It can be observed that if the system is uncontrolled, some resources are under-utilized while others are over-utilized.
@@
We work with $N=100$ players and $d=5$, and action sets $\{\boldsymbol{x}_{n}\mid\sum_{i=1}^{d}x_{n}^{i}\leq 10\}$ for each player $n$. Recall that the global cost is given by $\boldsymbol{x}^{\top}\boldsymbol{G}\boldsymbol{x}+\boldsymbol{\rho}^{T}\boldsymbol{x}$. To generate $\boldsymbol{G}$, we first generate random $\boldsymbol{H_{0}}\in\mathbb{R}^{Nd\times(Nd-5)}$ with each element of $\boldsymbol{H_{0}}$ sampled independently from a standard Gaussian distribution. We use $\boldsymbol{G}=\boldsymbol{H_{0}}\boldsymbol{H_{0}}^{T}$. A random $\boldsymbol{\rho}$ is obtained with each element sampled from a Gaussian distribution with mean 0 and variance $0.5$. The above matrices $\boldsymbol{Q_{n}}$ are obtained by sampling $\boldsymbol{H_{n}}\in\mathbb{R}^{Nd\times Nd}$ and using $\boldsymbol{Q_{n}}=\boldsymbol{H_{n}}\boldsymbol{H_{n}}^{T}$. The vectors $\boldsymbol{c_{n}}$ are sampled in a similar way as $\boldsymbol{\rho}$.